John, twelve years old, is three times as old as his brother. How old will John be when he is twice as old as his brother?
Answer by jrfrunner(365)

John, twelve years old, is three times as old as his brother. How old will John be when he is twice as old as his brother?
===
Let John's age = J = 12
Let John's brother's age = B
--
TODAY
J = 3*B is given
since J=12, then B=4
Today John's brother is 4 years old
---
future: X years away
J+X is John's age x years away
B+X is John's brother's age x years away
---
How old will John be when he is twice as old as his brother?
we want to find J+X
--
J+X=2*(B+X) we know today J=12, B=4 so substitute and solve for X
12+x=2*(4+X)
12+x=8+2x
4=x (four years in the future the above relationship will hold true
so J+X = 12+4=16
When John is 16 he will be twice his brother's age (B+X) = 4+4=8